CH/Agen/93/01

Agenda

Charissa Group Meeting

To be held on: 10 February 1993, 11.15 am

To be held at: Birmingham

Distribution: WNC,NMC,RAC,BRF,GJG,SJAH,RAH,JSL,ACM, VP,WDMR,GT,RPW,DLW

1.  Minutes of Meeting of 25 November 1992
2.  Matters Arising/ Actions
3.  Update on NSC/SERC matters  (DLW)
4.  Progress on detector tests  (DLW)
5.  Progress on scintillators  (NMC)
6.  Progress on chamber & detector assemblies  (JSL)
7.  Progress on gas supply  (RAC)
8.  Progress on electronics & data acquisition  (RAH)
9.  Developments on hybrid detectors  (RAC)
10.  Developments on event manager & GEC emulator  (VP)
11.  Documentation of Experimental Setup  (BRF)
12.  Completion of Programme at NSF  (BRF)
13.  ANU Planning  (BRF)

(1) Progress so far (plans and moving of equipment)

(2) Dates for running, and who covers what dates (3) Subsistence costs etc.

14.  Plans for Strasbourg
15. Legnaro Meeting (DLW,JSL)
16.  Minor Equipment Bids
17.  Glasgow Contributions
18.  Any other business
19.  Date of next meeting

W.N. Catford (Secretary) 9 February 1992